OpenStudy (anonymous):

What is the equation of the line that passes through (1, 2) and is parallel to the line whose equation is 4x + y - 1 = 0? 4x - y - 6 = 0 or 4x + y - 6 = 0 or 4x + y + 6 = 0

OpenStudy (anonymous):

Take the equation you're given and get it in y= form: \[y=-4x+1\] The slope on this line is -4. To be parallel, two lines must have the same slope, so the answer will have -4 as the slope. We have a point on the line and the slope so we can plug into this formula: \[y-y_1=m(x-x_1)\] so: \[y-2=-4(x-1)\] Simplify this and we get our answer: \[y=-4x+6\] And you can rearrange your answer choices to find the equivalent one.

OpenStudy (anonymous):

it would be the first one right

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

u subtract y from both sides right

OpenStudy (anonymous):

It would be the second one, subtract 4x and add 6 to both sides.

OpenStudy (anonymous):

yea i had that the 1st time

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're welcome, glad I could help.